### v1

# An approach to non-equilibrium statistical physics using variational Bayesian inference

Abstract
We discuss an approach to mathematically modelling systems made of objects that are coupled together, using generative models of the dependence relationships between states (or trajectories) of the things comprising such systems. This broad class includes open or non-equilibrium systems and is especially relevant to self-organising systems. The ensuing variational free energy principle (FEP) has certain advantages over using random dynamical systems explicitly, notably, by being more tractable and offering a parsimonious explanation of why the joint system evolves in the way that it does, based on the properties of the coupling between system components. Using the FEP allows us to model the dynamics of an object as if it were a process of variational inference, because variational free energy (or surprisal) is a Lyapunov function for its dynamics. In short, we argue that using generative models to represent and track relations among subsystems leads us to a particular statistical theory of interacting systems. Conversely, this theory enables us to construct nested models that respect the known relations among subsystems. We point out that the fact that a physical object conforms to the FEP does not necessarily imply that this object performs inference in the literal sense; rather, it is a useful explanatory fiction which replaces the ‘explicit’ dynamics of the object with an ‘implicit’ flow on free energy gradients—a fiction that may or may not be entertained by the object itself.
Acknowledgements
This paper more fully develops the theory outlined in a previous manuscript [1], and we are grateful to those listed there for their contributions to preliminary discussions on the topic. The account here is largely based on a series of talks given by the second named author at the Department of Philosophy of the University of Vienna and the Laufer Centre for Physical and Quantitative Biology, and the informal discussions that followed.
1 Introduction
In statistics, generative models are joint probability distributions that model the relationships between different variables—most often, observations and their causes, or data and a set of labels. In the study of open and interacting systems, the use of statistical models—capable of describing such relationships—seems to be particularly apt. One approach to making such models, using variational (approximate Bayesian) inference [2, 3], describes how a physical object reflects the properties of its environment and vice versa. Bayesian mechanics under the free energy principle (FEP) [4, 5] aims to describe the physical properties of a system in terms of dynamics on a manifold of probability distributions (i.e., as variational inference), resting on the useful property that physical objects generally look as if they track or infer the statistics of things to which they are coupled.
In this paper, we argue that generative models provide an especially useful foundation in scientific modelling generally, and especially in the study of systems exhibiting adaptation, morphogenesis, and other non-equilibrium phenomena. More specifically, we argue that generative modelling—premised on the FEP—is particularly apt for the practice of modelling the activity of modelling itself; for instance, in modelling sophisticated self-organising systems, where we write down a representation of the model that we assume is in play in prediction and action. We use the FEP to illustrate how implementing models which (i) formalise how one object keeps track of another, and (ii) codify these relationships, can increase the fidelity of our mathematical modelling. We identify a particular issue, which the FEP is able to uniquely address: the FEP keeps track of how a model depends on aspects of the thing that is modelled in a particularly insightful way.
Interestingly, because it incorporates this information about things that are coupled and their coupling, it has been suggested that the FEP conflates the metaphorical map (our mathematical model of the thing modelled) and the territory (the thing modelled). The so-called model reification or map-territory fallacy is a general critique of the practice of scientific modelling—in physics, principally due to Cartwright [6, 7]. Some have referenced this fallacy in relation to the FEP [8, 9]. In particular, it has been suggested that FEP-theoretic modelling conflates the metaphorical “map”—(i.e., the scientific model that scientists use to make sense of some real-world phenomenon)—and “territory,” i.e., the actual physical system that is being modelled. (See [10, 11, 12, 3] for related critical discussions of such claims.)
According to this line of thinking, using Bayesian mechanics to claim that we can model objects as themselves engaging in inference about the statistics of their environment constitutes a case of model reification. The allegation is that, in describing the dynamics of physical objects as implementing a form of inference—as opposed to considering the inferential aspects of the explanation as pertaining to our scientific models of those objects—the FEP theorist mistakenly conflates their metaphorical “map” of the territory (i.e., the scientific FEP-theoretic or Bayesian mechanical model) with the “territory” (the real-world system itself). This constitutes a reification of our model by assuming some aspect of the model or “map” is a real feature of the physical world or “territory” [10, 11]. By clarifying that this potential difficulty does not directly and necessarily apply to FEP-theoretic modelling, we are able to turn the story on its head: namely, we point out that the FEP provides a set of ultimate explanatory constraints on what counts as a model (or map) of anything whatsoever that exists in the physical world.
Here, we discuss the use of the FEP in statistics and statistical physics, and what advantages it offers; we then argue that there is no conflation of map and territory, i.e., of modelled system and scientific model, in the FEP—in fact, we describe how they are carefully distinguished. We argue that modelling the dynamics of an object such that it looks as if it models the system it is coupled to is not an instance of the map-territory fallacy. This argument gives way to a more general theme of the FEP, which is that capturing interactions and dependence relations between things using generative models is insightful. We discuss how using the FEP gives us a flexible model of the properties of the object that look like they perform inference, but crucially, without inherently reifying the inferential aspects of the model, attributing them to the real-world object. That is to say, a model of an object can always be written as modelling or inferring the structure of its environment. To make this argument, we appeal to two distinctions: one between the generative model (which is our scientific model) and the physical object being modelled (the real ‘thing’ in the world) [3]; and the other between the generative model and the variational density, which are often conflated in the secondary literature [13, 14].
To sketch the argument: the FEP provides us with tools from dynamical systems and information theory, allowing us to model any coupled random dynamical systems as statistical estimators, i.e., as engaging in the statistical estimation of some quantity pertaining to the set of objects with which it interacts. Under the FEP, one object that is coupled to any other can thus be read as a representation of the coupled or estimated object, albeit of a deflationary, qualified sort (discussed in [15]; see also [13, 16, 3]). More formally, the FEP applies to generative models that contain a partition of the states or paths of a joint system (a so called “particular” partition, i.e., into objects or particles). This partition induces a statistical boundary called a Markov blanket [17], which captures conditional independence relations amongst subsets of the system [2, 18]. The FEP says that, when such Markov blanketed subsets of the system exist, those subsystems track [19] each others’ statistical properties (dynamically) across the blanket. Essentially, the FEP is a modelling framework which captures the following fact: models of coupled particles or factorised systems can be written as parameter estimators, and estimating the parameters of a distribution is (a form of) Bayesian inference. We can think of the FEP, metaphorically, as a map of the part of the territory that behaves as if it were a map—or looks as such to an external observer.
Accounts such as [10] have encouraged practitioners of the FEP to be explicit about this point. Viewing the theory this way, not only is there no conflation between our FEP-theoretic model of the system and the system itself, but in fact we arrive at a novel modelling method for coupled random dynamical systems. In particular, we do not claim that the objects in the system themselves must literally be performing inference. Rather, we describe the dynamics as doing inference, in the sense of being a statistical estimator whose states parameterise a variational density.
We first present a conceptual, and then technical, introduction to the FEP and Bayesian mechanics—and the way it leverages the relationships encoded by a generative model to build a scientific model of the dynamics of physical objects using free energy gradients. This is possible by modelling the property of physical systems that they look as if they engage in modelling to an external observer. We then argue that there is no ambiguity between model and modelled real-world system in this formalism. We will examine possible reasons for mistakenly invoking the map-territory fallacy in this context, due mainly to the historically botched distinctions between the generative model and the variational density, on the one hand, and between the generative model and the so-called generative process, on the other.
We propose that the FEP and Bayesian mechanics provide a map of any possible map whatsoever—of, or held by—a physical system. (This includes the use of the FEP itself by a modeller.) We suggest that this leads, in a Wittgensteinian manner, to constraints on all possible maps, or models of physical things. This is meant in terms of logically necessary preconditions for sense-making, and thus the parsing of sensory streams on which cognition is contingent [15]. In particular, these constraints arise from what it means to be a map or model of a physical process at all.
We remark that we will not claim to invalidate the map-territory fallacy itself. That is, we do not suggest that the map and the territory are never conflated; they obviously are in cases of genuine model reification (as discussed cogently by [11]). Models built using the FEP are still scientific models, and the properties of FEP theoretic model also run the risk of reification. Instead, we will address the claim that the FEP is limited in its effectiveness or engages in model reification because real physical things cannot be modelled as themselves modelling their environments by minimising free energy.
2 The free energy principle
2.1 Conceptually
In this subsection we will provide an up-to-date overview of the concepts behind the FEP. One can also see [3] for a discussion of the main ideas. See the following subsection for more mathematical details.
Conceptually, the FEP starts from a partitions an overall “system” into three component sets of states: the internal states $\mu$ of a ‘thing’ or ‘particle’, an environment (which possesses external states $\eta$, relative to the thing in question), and an interface separating but coupling the two (the so-called Markov blanket, $b$), which mediates the influence of $\eta$ on $\mu$ and vice-versa. The states of the system will be tuples $x\coloneqq(\eta,b,\mu)$, where $\eta$ and $\mu$ are states of two subsystems and $b$ is some locus of states such that $p(\eta\mid\mu,b)=p(\eta\mid b)$ and vice versa. “Particular states” are the internal states of each thing and their respective Markov blanket. Of note, the particular partition is akin to choosing a set of coarse-grained variables at a superordinate scale, associating each state or path at the subordinate scale to a particle at the superordinate scale (as either an internal or a blanket state), in terms of the relations of conditional independence that obtain between states or paths [20, 2].
More precisely, we will consider systems of interacting objects, such as an open random dynamical system coupled to its environment. Fixing a final time $\tau$, the joint system will be denoted by the time-indexed family of random variables $\{X\}_{t:\tau}$ taking values $x\in\mathcal{X}$ at each $t$, with a specific realisation of the system (meaning, a sample path or trajectory beginning at some fixed $X_{0}=x_{0}$) being the sequence $\gamma(x_{0})=\{x_{0},x_{1},\ldots,x_{\tau}\}$. We will define $\gamma(x_{0})\mapsto\gamma_{t}(x_{0})$ as the projection to the state at time $t$. Then $B_{t}$ is what we have called the Markov blanket in $\{X\}_{t:\tau}$ at time $t$, separating it into statistically distinct subsystems (which one can then argue do not mix into each other [18]). We will repeat this notation for
$\gamma_{t}(\eta_{0})=N_{t}=\eta\in\mathcal{N}$
$\gamma_{t}(b_{0})=B_{t}=b\in\mathcal{B}$
$\gamma_{t}(\mu_{0})=M_{t}=\mu\in\mathcal{M}$
as well as for conditional trajectories such as
$\gamma_{t}(\eta_{0}\mid B_{t}=b)=\eta_{b}\in\mathcal{N}\crossproduct\{b\}.$
Although we will not go into the theory of random dynamical systems, when equipped with a time evolution map acting on the sample space, these are all the ingredients necessary to define a random dynamical system [21], and this is implicit in our discussion; for details see the account in [14]. When $\{M\}_{t:\tau}$ is an open system coupled to an environment with states $\eta$, a value $\mu$ is called an internal state, and the subsystem $\{M\}_{t:\tau}$ is variously called an object, a thing, or a particle in the literature. This is the convention we will stick to.
We can guarantee that the dynamics of a system can be described as a flow on surprisal, and thereby, cast as implementing a form of inference, whenever there exists a “particular partition” of that system. At the core of the FEP is a mathematical statement about the behaviour of such systems. For a comprehensive overview, see [3]. In particular, the FEP allows us relate the dynamics of states or paths that are internal to the object, $\mu$, to the dynamics of external states, conditioned on boundary or blanket states $p(\eta\mid b)$—that is, understanding what goes on inside a subsystem by understanding what it is coupled to, as ‘seen’ via blanket states [5, 22].
The logic of the FEP can be summarised as follows. If there exists a particular partition of a system, then there exists a coupling between internal and external states or paths, given blanket states or paths; then, since the internal dynamics of a given object are coupled to other (external) subsets of the system—in such a way that they can be read as parameters for probability densities over the external subset—the FEP describes these interacting systems using variational inference. This is a core aspect of the FEP that distinguishes it from a straightforward application of density dynamics to the internal states per se.
This is where the interpretation of dynamics as a form of inference comes into play. Indeed, one way to state the FEP is:
Claim 1.
If a random dynamical system (i) is subject to a specific kind of sparse coupling or boundary condition, such that it possesses a Markov blanket, and (ii) has a subsystem $\{M\}_{t:\tau}$ under that blanket with characteristic states (in the sense that it occupies a pullback attractor), then $\{M\}_{t:\tau}$ can always be described as if it minimises a variational free energy functional of densities over states or trajectories of $\{N\}_{t:\tau}$ [14, 22].
Another statement one can make is that Bayesian mechanics formalises the truism that, on average, an object will do whatever it is that it does on average, given the sort of object it is. Behind this tautological statement, we find a formal apparatus that enables us to model a variety of dynamical systems in a parsimonious manner. We hope this will allow us to take some first steps towards a mechanistic description of how self-organisation manifests in physical systems with such pullback attractors [3].
Bayesian mechanics explains why every physical thing looks as if it were performing inference over the causes of its perturbations, forming (Bayes) optimal beliefs about those causes [22]. Bayesian mechanics begins from the minimisation of self-information (also called surprisal) or free energy, and allows us to write down mechanical theories (i.e., laws of motion) yielding the dynamics of systems on an attractor [19]. Surprisal is simply the negative log-probability of an observed outcome. To say that an object minimises (or follows a path of least) surprisal is to say that an object evolves towards the outcomes which are most probable, given the sort of thing that it is (meaning, its characteristic states of being, or characteristic trajectories through its state space).
Interestingly, this means that surprisal is suited to act as a kind of ‘ontological potential’ for the object, in the sense that the surprisal is equivalent to a constraint on what states and outcomes are accessible to the object; the trajectories of a physical system are random curves through the gradient of surprisal over the state space, which satisfy an equation of motion derived from the surprisal itself [4, 23]. As has been asserted in recent literature [19], one reading of the FEP would hold that the crux of the ubiquity of the FEP is that everything which is coupled to another thing holds a model of that thing to which it is coupled in the following, trivial sense: in virtue of this coupling, one subsystem estimates the parameters of the probability density over states (or trajectories) of the other subsystems (and vice versa). Using this property, they can be written as performing inference in the sense of variational Bayesian inference.
This coupling is sometimes discussed in terms of the ‘probabilistic beliefs’ that one object holds about its environment. Given that the term ‘belief’ is quite theoretically loaded (especially in philosophy), we ought perhaps to say ‘estimator,’ instead. Once more, our contention is that anything which estimates the parameters of a density is engaged in a form of Bayesian inference, by mathematical definition; i.e., it is mathematically indisputable that systems that estimate statistics represent, or encode, or otherwise hold probabilistic beliefs—whether the practitioner likes the word ‘beliefs’ or not, and irrespective of what commitments one has with respect to notions such as encoding and representation (e.g., [24]).
Generative models and variational methods allow us to licence these claims [25, 26]. Generative models are joint probability densities over causes and consequences which, here, are used to capture the dependencies between the variables constituting $\{X\}_{t:\tau}$, meaning the way that changes in some variables cause changes in others. Writing down a generative model for some real object in the world (the metaphorical territory) is building a metaphorical map of that territory (a formalisation or codification of the relations between the things in that territory).
Variational techniques comprise a family of inference techniques which allow a modeller to approximate some (usually intractable) posterior probability density under a (usually specified) generative model. In variational inference, we write down a parameterised probability density function called a variational density, which is used as an approximation to the posterior probability density. The quantity called variational free energy scores the difference between the variational density and the posterior over external states—here, a conditional density, given a particular blanket state. Hence, when some notion of the difference between them is zero, variational free energy can be used as a proxy for surprise, and we can approximate the posterior density; which is a form of parameterised Bayesian inference.
Note that any given object cannot change the generative model associated to that object. This follows because the object can only change the variational density encoded by its internal states. However, to explain these changes—under a variational principle of least action—one needs to specify the generative model under which the posterior is approximated. This means the generative model explains the dynamics of something that is observed in a particular way. This is usually expressed by saying that the generative model is entailed by the thing in question. That an object entails a generative model is consistent with statements like “an object is a model of its environment”: e.g., in the sense of the good regulator theorem [19]). This follows because the generative model just is a probabilistic description of the joint states an object and its environment occupy.
2.2 Mathematically
With that conceptual overview being stated, we will now provide some further mathematical details. Similar (contemporary) reviews can be found in [4, 5, 19, 27]. We will argue that coupled random dynamical systems partitioned under a Markov blanket naturally have the form of variational Bayesian inference problems, in the sense that we obtain the values of parameters from the minimisation of free energy or surprisal. This is what licences the claim that minimising free energy implements a form of inference, in which case the object models its environment. Ultimately this simply means that, in virtue of being coupled via blanket states to its environment, the object reflects the structure of, and data about (i.e., models), its environment—or, what it ‘believes’ that its environment is like (in the sense of being a statistical estimator, as defined above).
Appealing to the Laplace approximation (that if a density has a unique maximum, in a neighbourhood of the most likely state, then the density is approximately Gaussian with that mean), we will suppose all random variables and all trajectories are jointly Gaussian. We will call the conditionally expected state—that is, the expected state given some fixed blanket state—the conditional mode $\hat{\mu}_{b}$ or $\hat{\eta}_{b}$. Let $\mathcal{L}(\mu,b)=-\log p(\mu,b)$ be the surprisal of internal and blanket states, such that $\hat{\mu}_{b}=\operatorname*{\mathrm{arg\,min}}_{\mu}\mathcal{L}(\mu,b)$ and thus satisfies $\nabla_{\mu}\mathcal{L}(\mu,b)=0$. By the Freidlin–Wentzell theorem, we can argue the same is true of paths (see [19]). In that setting, the interpretation of surprisal as the fluctuations of a random dynamical system away from its expected path plays a crucial role in stochastic thermodynamics [28, 29].
The FEP states that, given the existence of a boundary, we can construct a map relating (the dynamics of) internal states to (those of) the environment across the boundary. In particular, if there exists a boundary in the form of a Markov blanket, then there exists a function $\sigma:\mathcal{M}\to\mathcal{N}$ such that $\sigma(\hat{\mu}_{b})=\hat{\eta}_{b}$, sending conditional modes to conditional modes [22, Lemma 4.3]. We construct this map by composing the natural projections to boundary states, $\bm{\mu}:\mathcal{B}\to\mathcal{M}$ and $\bm{\eta}:\mathcal{B}\to\mathcal{N}$, into
$\bm{\eta}\circ\bm{\mu}^{-1}:\mathcal{M}\to\mathcal{N}.$
We have assumed $\bm{\mu}$ is at minimum an injection, hence bijective on its image, from which we can take a left-cancellative pseudo-inverse. Under hypotheses on these maps, for example, that two blanket states which map to the same value of $\mu$ also map to the same value of $\eta$, this is a useful function, in the sense that it associates a unique pair of conditional modes to each blanket state. [^1] One can see [4] for a proof looking at the properties of solutions to certain linear systems of equations. Alternatively, note that this is evident from taking a tensor-Hom adjunct of the composition, $b\mapsto\xi(b,-)$, and asking that this map be an injection in the sense that when $i=j$, the map takes $\hat{\mu}_{i}=\hat{\mu}_{j}$ to equal images $\hat{\eta}_{i}=\hat{\eta}_{j}$ [22, pp 25]. Importantly, in the definition of $\sigma$, we can pass an arbitrary internal state in; its image will simply not be the conditional mode of the environment, and hence it will not parameterise a density consistent with $p(\eta\mid b)$.
We then postulate the existence of a family of probability densities parameterised by values of some known quantity, called the variational density $q(\eta;\hat{\eta})$. Generically this is constructed by taking a family of Gaussian densities with different modes. By the previous discussion, we can interpret this as the likely configuration of the environment, knowing the environment must be amenable to a certain internal state on average (given a corresponding blanket state), $q(\eta;\sigma(\mu_{b}))$. It is immediate that the probability density parameterised by the conditional mode $\hat{\eta}_{b}$ ought to be the same as the conditional probability $p(\eta\mid b)$. Moreover, if we assume the system is coupled by a particular partition, we naturally have $\sigma(\hat{\mu}_{b})=\hat{\eta}_{b}$. Together, we have supposed that the object exists as the preimage of $\hat{\eta}_{b}$, or tends to do so, and hence, that the environmental states that are likely under a given conditional mode of the object ought to be equivalent to the probabilities over external states conditioned on that blanket state.
Because we are identifying a parameter value which both (i) has the greatest log-likelihood of being observed and (ii) fits a parameterised distribution over unobserved states to a distribution over unobserved states given observed states, this set up naturally has the structure of a variational inference problem [30]. As discussed, the quantity known as variational free energy scores the difference between the variational density and the true conditional density in terms of the relative entropy from one to the other, known as the KL divergence, added to the surprisal of internal and blanket states:
[…]
In the context of variational inference, we consider both terms in order to obtain a parameter for which $q=p$ almost surely and identify the least surprising (or most likely) parameter; the interpretation of the surprisal term is obvious, whilst the entropy term prevents $q$ from concentrating about the mode, spreading to match $p$ [^2] Note that as before, we can tell this story for the trajectories of a subsystem, where we take $q(\gamma(\eta_{0});u_{b})$ and $p(\gamma(\eta_{0})\mid\gamma(b_{0}))$ instead, where $u_{b}$ is the expected path given an observed trajectory of blanket states. The use of maximum entropy in the setting of sample paths is reviewed in [31]. Again, in the context of variational inference, the least surprising parameter is also the parameter with the most evidence—meaning objects that are ‘good’ at occupying their characteristic states can be described as self-evidencing. This has the interpretation of minimising the surprisal of observations of $\mu,b$ by inferring probabilities of unobserved latent states $\eta$ using the so-called discriminative model $p(\eta\mid b,\mu)$.
Importantly, the variational free energy is always greater than or equal to the surprisal, since the KL divergence is non-negative in general. Minimising the free energy effectively implements a form of inference, where we have inferred the best possible approximation to the true conditional or posterior density. That is, when the KL term is zero, the variational density and the density over external states given the state of the blanket are equal up to sets of zero probability. We observe (by properties of $\sigma$) that this is only possible when $\mu=\hat{\mu}_{b}$. In the limit of exact Bayesian inference, when the functional form of posterior beliefs encoded by internal states coincides with the posterior density over external states, the variational free energy reduces to the self-information of internal and blanket states. (Refer to e.g. [22, Lemma 4.2], or the variational free energy lemma in [27] for more discussion.)
This returns us to our earlier discussion about parameter estimation in the FEP. In this case (and again invoking the Laplace approximation), the surprisal being non-zero has a neat interpretation as the uncertainty that the object cannot resolve by being at the mode. Since the object is conditionally independent of its environment, it cannot ‘see’ beyond the Markov blanket—$N_{t}$ is an unobserved latent variable causing observed blanket states—lending this inference process a physical interpretation appropriate for self-organising and adaptive objects as well.
We will now give the following result: the object is at its most likely point in state or path space if and only if it minimises variational free energy. We will recall that the KL divergence is non-negative, implying
$D_{\mathrm{KL}}(q,p)+\mathcal{L}(\mu,b)\geq\mathcal{L}(\mu,b)$
saturating at $\mu=\hat{\mu}_{b}$ by construction. Since we defined $\hat{\mu}_{b}=\operatorname*{\mathrm{arg\,min}}_{\mu}\mathcal{L}(\mu,b)$, and any movement in parameter space $\hat{\mu}_{b}+v$ adds a non-zero KL term to $\mathcal{L}(\hat{\mu}_{b},b)$, we conclude
$\nabla_{\mu}\mathcal{L}(\mu,b)=0\iff\nabla_{\mu}F(\mu,b)=0.$
What we have shown is there is a natural interpretation of the most likely state as a parameter estimation, in the sense of Bayesian inference. This argument, whilst logically straightforward, conceals a more subtle point: all of this begins from the existence of a generative model. In particular, since the joint probability $p(\eta,b,\mu)$ factorises into $p(\eta\mid b,\mu)p(b,\mu)$ (see the probabilistic chain rule), logarithms of products decompose additively, and we have $p(\eta\mid b,\mu)=p(\eta\mid b)$ by assumption, we have
[…]
In particular, if we begin with a generative model and a variational density, we are led to an inference with a satisfactory physical interpretation and explanatory power (we will dive into the latter in the section on dynamics). Not only is this the case, but by elementary mathematical considerations [^3] Notice that the support of the generative model and variational density differ. Whilst we may still take the KL divergence if the variational density is absolutely continuous with respect to the generative model, this should not be expected to be meaningful. we find we ought to take this factorisation. This observation is the main point of our proposal. The use of this form of the free energy has an even more straightforward interpretation as using an evidence lower bound to infer the generative model, such that the object can generate samples of how environmental states are likely to vary with internal states. Our objective in using the FEP is to use this mathematical fact to replace classic models of stochastic processes—which use the minimisation of surprisal—with a more tractable and interpretable quantity: the variational free energy.
2.2.1 An overview so far
All of this suggests that if a model of a subsystem occupies the states which we expect it to, given what the subsystem is, then we can model that subsystem as performing inference (see [19]). The idea is predicated on the tautology that so long as we observe that subsystem as being ‘the thing that it is’, then it must (by definition) be in a state that is amenable to its continued existence as that kind of thing. Similarly, if an object remains identifiable as a particular thing through time, it must be evolving in the way that we expect of that thing. Aaronson refers to this idea—the existence of a system in a certain state being conditioned on the result of a computation (and therefore encoding that computation in its state)—as “anthropic computation” [32]. One phrasing of our proposal is that using generative models in the study of coupled random dynamical systems is a way of leveraging anthropic computation to build models of systems in statistical physics, and that the FEP gives a formal apparatus with which we can build such models.
Put a different way, the FEP provides us with tools that allow us to model any particle, or “thing”—in the sense that it exists as a separable, identifiable thing—as if it were inferring the external states causing its boundary states, and (re)acting so as to remain in system-like configurations. This interpretation of the dynamics as inference comes from the minimisation of a free energy functional of beliefs about the causes of blanket states, as parameterised by internal states, which guarantees—or describes—the minimisation of the surprisal of $(\mu,b)$. Using this approach, we can understand the dynamics or observable behaviour of things as realising a path of minimal free energy [3, 27]. Since the object cannot ‘see’ past its Markov blanket, it must infer the external states it is interacting with; moreover, when it does so, it can be said to maximise self-evidence.
It is interesting to note that there is no guarantee the system’s model is good; in fact, we expect simple systems with poor ability to self-sustain to have simple models, meaning the FEP gives us concrete predictions about how “anthropic computers” might behave when it is too difficult to minimise a given potential. This has been discussed in the context of the FEP in [33, 34].
2.3 Dynamically
We will now explore what the FEP says about the dynamics of objects that minimise variational free energy, focusing now on the writing of a dynamical system in terms of free energy gradients—and how it relates physical properties of an object to the dynamics of beliefs parameterised by the likely states or trajectories of that object. For similar reviews one can see [2, 4, 19, 27]. We will give a quick example of how the heat dissipated into the environment by an object behind a boundary maintaining a non-equilibrium steady state can be written in terms of the variational free energy.
Assume the state space $\mathcal{N}$ is some open, flat Riemannian manifold of dimension $n$ ($\dim\mathcal{B}=d$, $\dim\mathcal{M}=m$, respectively), where we set $\ell=n+d+m$. Let $H_{t}$ be a one-dimensional standard Brownian motion: $\{H_{t}\}_{t:\tau}$ has stationary and independent increments, $H_{0}=0$ almost surely, and $t\mapsto H_{t}$ is continuous almost surely for all $t\in[0,T]$. Now let $W_{t}$ be a $k$-vector of independent Brownian motions (also called a $k$-dimensional Wiener process).
By elementary arguments in stochastic analysis, we know that we can write any stochastic differential equation (SDE) with an equilibrium measure as
[…]
Arguments in the literature imply this holds for non-equilibrium steady states as well, with the matrices described in [35, 36]:
[…]
where $\nu$ is an $\ell$-by-$k$ Lipshitz continuous matrix field with bounded linear growth. In stochastic thermodynamics the surprisal of a path given a particular ensemble of paths is often used—in our notation,
[…]
where the entries in the diffusion matrix may change in time. In each situation we have a vector field—the gradient flow of surprisal—added to a Gaussian distributed random variable with mean zero and variance $t$. (Note that from here onwards we will fix all fluctuation- and curvature-related constants to $+1$.)
Recall that $\nabla_{\mu}\mathcal{L}(\mu,b)=0$ if and only if $\nabla_{\mu}F(\mu,b)=0$. An infinitesimal variation in the KL divergence is the Fisher information; hence, using the Laplace approximation, in a neighbourhood of the mode the KL term is a constant. In that case, for a given blanket state $B_{t}=\beta$ (fixed in time) the conditional mode evolves [^4] With error bounded locally by some cubic polynomial in $(\mu-\hat{\mu}_{\beta})$. Note that since $q_{\mu}(\eta)$ is exponential by assumption, the Hessian of this free energy is non-singular and positive-definite, such that (i) second-order terms suffice to describe its behaviour in a neighbourhood of the minimum and (ii) first-order terms vanish at the minimum. as
$\partial_{t}\hat{\mu}_{\beta}(t)=-\nabla_{\mu}F(\mu,\beta)$
coming from the following SDE on the state space ××$\mathcal{M}\crossproduct\mathcal{B}\crossproduct\mathbf{R}^{2k}$:
[…]
and so on for the other equations above.
Here we recover our original notion of replacing the dynamics of the object with an inference process: the dynamics of the most likely state of the object approximately perform a minimisation of the variational free energy functional. In the path setting this becomes even more apparent, since the most likely path simply is the solution to the equations of motion minimising free energy.
2.3.1 Stochastic thermodynamics
We may now ask whether the dynamics of beliefs, seen as performing variational inference, can be related to the physical dynamics of states parameterising those beliefs. We show that we can relate performing inference to dissipating heat using this FEP.
We will consider a thermostatted system—an object embedded in a heat bath of temperature $T$ called a medium or an environment—at non-equilibrium steady state, such as a colloidal particle moving in one spatial dimension. The force on the particle is a function of both its state and a time-independent control parameter [28]. We set the blanket state as the control parameter, since this determines the mode of the non-equilibrium steady state. In this case, since we have set $\partial_{t}\hat{\mu}_{b}(t)=-\nabla F(\mu,b)$ in a neighbourhood of the mode, under mild estimates on the expectation we have the Langevin equation
$\partial_{t}\mu(t)=-\nabla_{\mu}F(\mu_{t},b)+\xi_{t}$
whose solutions for different realisations of noise are trajectories of the particle, $\gamma(\mu_{0})$. As in Freidlin–Wentzell theory, such trajectories are generated with probability
[…]
which is simply the magnitude of fluctuations (with the extra term motivated in [37]).
[…]
$\mathbf{q}[\gamma(\mu_{0})]=T\log\frac{p(\gamma(\mu_{0}))}{p(\tilde{\gamma}(\mu_{0}))}$ (1)
where $\tilde{\gamma}(\mu_{0})$ is a time-reversed path.
Computing the difference contained in (1), the heat dissipated into the environment is
[…]
This construction could be read as an implicit thermodynamics of belief updating; in the sense that it expresses the work done by a system in terms of inference. This nicely recapitulates the formulation of thermodynamics in terms of information length, based upon dynamics (of internal states) [40, 41]. See also: [4]. For example, [40] concludes that “Bennett’s classic acceptance ratio method for measuring free energy differences also measures thermodynamic length”. However, the free energy differences in this setting refer to thermodynamic free energy differences of internal states. The dénouement of the current treatment is that the same conclusions apply to variational free energy and, implicitly, beliefs about external states held by internal states.
3 Summary remarks
To summarise: formulated most generally, the FEP says that the average path taken by an object that is sparsely coupled to its environment is characterised by a Lyapunov function—a function that decreases monotonically along the trajectories of a (deterministic) dynamical system. The utility of Lyapunov functions is that they guarantee stability, allowing us to understand the dynamical system as flowing on the gradient of the Lyapunov function. In the FEP, this Lyapunov function is surprisal or variational free energy, depending on the setting. In other words, to understand the claim that the FEP models a system as doing inference, it is crucial to understanding the ‘inference’ component of surprisal minimisation. Surprisal is simply the negative log probability. Variational free energy is a tractable (i.e., easily computed) upper bound on surprisal, where minimising variational free energy allows us to infer a target density by approximating it with a variational density. The surprisal being considered is, much like probability, surprisal given a model, in that it quantifies how surprising it would be for an observer—equipped with a model—to find the system (such as itself) in some non-system-like state or configuration. Minimising surprisal implies that the variational free energy is minimised (i.e. the object-as-model is a good estimator of its environment [19]), and likewise, minimising the variational free energy places an upper bound on the surprisal.
The crucial idea is that the dynamics of a system does inference when written as minimising a free energy functional, which is itself the consequence of reading surprisal as a Lyapunov function for those dynamics. In turn, the dynamics of a system minimise surprisal or free energy whenever that system is equipped with a Markov blanket; and hence, the dynamics of a system with a particular partition can always be written as a gradient flow on surprisal. This gradient flow is what we identify as actually being the process (i.e., the performance) of inference. Explicit examples of this can be found in [23, 42].
More specifically, suppose there exists a conditional density over external states, $q(\eta;\hat{\eta}_{b},\vartheta_{\eta\mid b})$, where we have parameterised the variational density by the conditional mode $\hat{\eta}_{b}$ and variance $\vartheta_{\eta\mid b}$ of the environment. For convenience, we assume a Laplace approximation to the posterior, meaning that in a neighbourhood of its peak, the posterior is approximately Gaussian, and hence uniquely determined by its mean and variance. This guarantees that finding the right values of $\hat{\eta}_{b},\vartheta_{\eta\mid b}$ are sufficient for $q$ to equal the posterior almost surely. More generally, under the maximum entropy assumption of a Gibbs density, any constraint function suffices.
The FEP says that if there exists a boundary in the form of a Markov blanket, then there exists a function $\sigma$ relating (the dynamics of) internal states to the environment across the boundary [22, Lemma 4.3]. In that case, we can map these parameters to $\sigma(\hat{\mu}_{b})$ and $[\eta-\sigma(\hat{\mu}_{b})]^{2}$, meaning that the particle models its environment. Ultimately this simply means that, in virtue of being coupled (via blanket states) to its environment, the particle reflects data about its environment—or, what it believes that its environment is like. Under the assumption that a particle which exists in an environment models its environment, free energy is implicitly a measurement of coherence. If the particle stops modelling its environment in a particular way (i.e., for a particular $\sigma$), it must have ceased to exist with the mean that $\sigma$ was constructed for; conversely, if a particle ceases to exist, it will stop modelling its environment.
If free energy (or surprisal) remains irreducibly high, it is likely that the system occupies surprising (i.e., implausible or uncharacteristic) states; this would render it non-existent over some timescale, in the sense that it would be occupying states that were not characteristic of the thing in question (e.g., a fish out of water). These statements tie together the reasoning that things which exist in an environment reflect the statistics of their environment in a particular way, based on the way they are coupled to their environment; or, they do not exist that way.
Similarly, note the direction of the implication in Claim 1. In the context of modelling persistent systems, the FEP is a sort of inversion of the typical reasoning one sees: instead of saying an object persists (as a thing) because it is stable, we say an object is stable (as what it is) because it persists (as that thing) [22]. We stipulate that the KL divergence vanishes—because it is a tautology that $p$ should equal $q$ almost surely—and because physics only ‘works’ in a given regime if this is the case (see [19, §IIB] where this argument is spelt out). This happens in the background, so to speak. Then we can use the FEP to describe the dynamics as a gradient flow on surprisal—which can be done by a simple mathematical fact and independently of the FEP—as a process which has performed inference in virtue of existing. Adding this ‘inflationary’ KL term, the non-equilibrium steady state density of the system can be understood as a model of the environment under this stipulation. All this is made possible by the existence of a particular partition. Hence, we arrive at a deflationary account of intelligence through an inflationary mathematical technique.
From the viewpoint of model building, since internal states are by construction unobservable, we can in principle tell any story that we like about them—provided that this story is supported well enough by observable evidence. One such story is that those internal states look as if they parameterise a variational density over external states. Taking this explanation seriously, and leveraging it as an approach to modelling physical systems, is applying the FEP.
4 Some preliminary philosophical considerations
We will now enter a more philosophical discussion concerning the use of the FEP. The preceding technical arguments establish the FEP as an explanatory principle that can be applied to any thing (particle, person, or population) around which one can draw a boundary or Markov blanket. This means that the FEP is what it says on the tin: it is a principle. In the physics, this means that the FEP is the foundation of a method to be applied [3]. So, how might one apply the FEP?
At the most general level of analysis, it could be argued that any scientific process whatsoever is an application of the FEP: in the sense that any scientific explanation or model involves resolving uncertainty about some observed thing or dynamics, under some hypotheses about the observed world.
More specifically, how might one apply the FEP to a particular particle, person or population? The answer is straightforward: the observer is trying to infer the generative model—which contains unobservable internal states of an object—that best explains the observed dynamics of some thing. What is observed is simply the blanket states of some ‘thing’ that shield internal states from direct observation. If a generative model that best explains an object’s behaviour can be identified, then by construction the FEP affords the modeller a complete (but not over-complete) description of the object’s internal dynamics, even though they can never be observed. As noted above, technical term for this relationship between real-world system and generative model is “entailment”: where a system that conforms to the dependence structure of the generative model is said to entail that generative model.
In short, if we equate the territory with the full joint system (crucially, including the unobserved internal states or dynamics of the observed thing), then the generative model plays the role of a map that best describes that (partially observed) territory. Note here, that the map only exists in relation to the observer—and, crucially, that the observer could be observing itself. In addition, applications of the FEP enables one to map out the parts of a territory that are unobservable in another sense; namely, in the sense that one cannot observe every state a given thing has been, or will be in. And even if one could, the internal states would be forever sequestered behind blanket states.
In short, applications of the FEP place the generative model centre stage as explanations for the behaviour of things—in terms of a map of some territory that can only be observed through its impressions on its Markov blanket.
5 A map of that part of the territory that behaves as if it were a map
We argue that the FEP can be understood, metaphorically, a “map” of sorts: a map of that part of the territory which behaves or looks as if it were a map. In other words, the FEP provides us with tools to understand the mathematically striking property of self-organising systems: that they look as if they infer, track, reflect, or represent the statistics of their environment. Even more simply, it gives us tools to model systems that look as if they are modelling the world.
How does FEP-theoretic technology allow us to model self-organising systems as modelling their world? The generative model or joint probability $p(\eta,b,\mu)$ plays a central role. This generative model can be read as a probabilistic specification of the states or paths characteristically occupied by the joint particle-environment system. In other words, our scientific model (or map) includes the generative model entailed by the system (i.e., the territory). In this sense, the generative model we hypothesise—to account for observations of an object—allows us to represent the constraints on the kinds of states the joint system can be found in. In other words, the generative model is a modeller’s map of the whole “territory” that characterises the observed system and its environment (including the modeller, as an observer in the environment).
The technology of the FEP allows us to say that certain components of the system being considered look as if they track other components, and that this is a feature of our map contained in a generative model under a particular partition. In providing us with a calculus allowing us to model things as modelling other things, the FEP provides us with a map (a scientific model) of any possible map or tracking relationship between things (namely, a mapping between internal and external states, mediated by blanket states). [^5] Here, we are assuming representationalist accounts of scientific practice. For alternatives, see [10].
Does our ability to always describe some dynamics as a form of inference imply that those dynamics are literally a form of inference, as opposed to being a manner of describing those dynamics? Here, it is important to defuse a possible (but in our view naïve) objection. It is a truism that physical systems need not explicitly calculate their trajectories of motion, to be modelled as pursuing such trajectories. In developing a Bayesian mechanical account of the dynamics of systems, we need not assume that the particle itself is literally performing inference. [^6] Ways of defending stronger, increasingly literal versions of this positions are available as well. For a defence of the claim that physical systems are quite literally in the business of inference, see [43, 44]. Related to this, views of physical systems implementing computations has been articulated [45, 32], as well as views where systems self-organise by computing predictions [46, 47, 42, 29]. What is at stake with the FEP is an ‘as if’ description. We simply take the fact that surprisal varies with the dynamics of the system, which implies free energy is minimised; the minimisation of this quantity is mathematically equivalent to inference, in the sense of being an estimator.
The fact that it is our model of the system which does inference and not necessarily the system itself is both a boon and a bane. It enables us to model the system using the more tractable gradient of variational free energy and the mathematics of statistical inference, rather than studying the possibly highly non-linear coupling between the dynamics of the two systems. However, it is not necessarily a faithful representation of the system in the literalist sense. See also [10].
In summary, our scientific, FEP-theoretic model is, metaphorically, a map that allows us to say that some (internal) subset of the system looks as if it possesses a map; which we interpret formally as tracking the statistics of another (external) subset. Our map (as scientists and modellers) can be identified in an unproblematic way as the generative model of the system, which FEP-theoretic technology enables us to write down. It is by construction that our model is a model of the modelling capacities of subsets of the real-world system: that is, our map is precisely a map of how objects behave as if they were maps, allowing us to construct scientific models of the maps encoded (or looking as if they were encoded) by the internal subset of the system considered. This is an answer to allegations that FEP-theoretic modelling commits the map-territory fallacy: namely, it shows that there is no conflation of the map and the territory, in the sense of the map we postulate as being embodied by the system within the mathematical machinery of the FEP.
Taking all the above, we might say that the FEP ultimately provides us with a map of any possible map whatsoever of, or held by, a physical system. This echoes seminal work in philosophy by Wittgenstein [48]. In his famous Tractatus Logico-Philosophicus, Wittgenstein set out to delimit the domain of sensible propositions by determining the general form of a proposition, an utterance able to carry a truth value. This set ultimate constraints on what is sensible, and bounded what is meaningful “from inside” of language itself. Similarly, the FEP provides ultimate constraints on possible maps or modelling relations, which arise from what it means to be a map or model of a physical process at all. Indeed, any modelling process consistent with the laws of physics must conform to the FEP, such that the FEP sets ultimate limits on what can count as a map or model—starting from within the technology of mapping or modelling itself.
In sum, the generative model is a kind of scientific model, which harnesses what we know about the dependency structure of the system. A variational density, on the other hand, is a probability density over external states or paths, implemented in terms of the sufficient statistics of external states given blanket states. It is an explicitly inferential representation of the manner in which subsets of a system with a particular partition track each other: in which case, we are describing a part of the territory that, as seen in our map, also behaves as if it were a map.
6 Some remarks on terminological idiosyncrasies of the FEP literature
In summary, we have argued that the map-territory fallacy, as it has been leveraged against the FEP (e.g., by [8, 9]) does not apply to FEP-theoretic modelling—that is, not in the sense that it conflates the model and the thing modelled. This allegation constitutes a fallacy, which one might call the map-territory fallacy fallacy.
Mathematically, there is no ambiguity or conflation of map and territory in FEP-theoretic modelling. Although the language used to narrate the mathematics of the FEP are at times ambiguous (which may have led to the confusion in philosophical commentary on the FEP), the mathematical structure itself is not. In the use of the FEP, an object is modelled as deploying a variational density, which its internal states parameterise; its dynamics are given in our generative model, which is meant to capture or represent the real (or most likely) dependence relations of the real-world system [13]. The It is thus a scientific principle, allowing us (as modellers) to construct scientific models (implemented as a generative model of states or paths), the content of which is precisely a mechanistic theory for the qualified representational capacities of certain particular systems (the property that they look as if they infer or track other things in the system, to which they are coupled) [3].
Charitably, one might argue that the allegations that FEP-theoretic modelling commits the map-territory fallacy may be due to several terminological ambiguities and idiosyncrasies that have mired the FEP literature. In fairness, these ambiguities and idiosyncrasies may lend themselves to misinterpretation. We now discuss two such cases.
6.1 Generative models and variational densities versus recognition densities
The first concerns the ambiguous discussion or description of two core constructs in the FEP literature—namely, the generative model and the variational (or recognition) density. It is noteworthy that the manner in which these constructs are deployed in the FEP-theoretic formulation is markedly different from previous Bayesian approaches. This has seldom been noted in the literature (with a few exceptions, e.g., [13]); and even when it is discussed, additional issues tend to undermine the clarity of the exposition (as noted by [9]).
In more traditional machine learning approaches, e.g., in the literature on the Helmholtz machine [49], the generative model and the recognition model correspond to the feedforward and the feedback passes that together constitute the process of data-driven Bayesian inference. In these settings, both the generative model and the recognition model are straightforwardly encoded in the connection weights of the Helmholtz machine: the generative model is encoded in the weighting of the backwards connections, which establish an information channel from high-level latent spaces of the neural network to its sensory layers; whereas the recognition model is encoded in the feedforward connections. In this setting, one assumes that any system that implements this pair of functionally differentiated weightings (e.g., a brain) implements a generative and recognition model pair.
Much of the literature on the FEP is based on this reading of the generative and recognition models: it is often said within this literature that FEP-theoretic generative models are encoded by the brain, much like these would be in a Helmholtz machine. However, whilst generative and variational models are also used in the FEP-theoretic formulation, they are deployed in a markedly different way, which we discussed extensively above.
Overlooking these differences may have contributed to the impression that theorists of the FEP conflate the metaphorical map and the territory. Recall that a generative model, generically, is just a joint probability density defined over the states (or paths) of a system. In other words, it is our map or representation of the system being modelled and its mechanics; and more particularly, it is our model of the dependencies between the variables of the system, providing us with equations of motion for the whole system (i.e., it is defined over internal, blanket, and external states/paths).
Technically, we say that the system that we are modelling entails a generative model; whereas the internal states of the particle (look, to the modeller, as if they) estimate the parameters of a conditional density over external states. This difference is crucial to understanding what is at stake in FEP-theoretic modelling: we say that one subset of variables over which the generative model is defined (i.e., the internal states) encodes a variational or recognition model over another subset (the external states). Entailment is the property that, metaphorically, “the part of the territory that behaves like a map” (the internal states that encode the parameters of the variational density) is itself part of the territory described by our map (the generative model) and so is subject to the mechanics of that territory, as described by our map. Entailment means that the relationships that hold between variables, as we represent in the generative model, are obtained as a description of the dynamics of the real-world system.
The subtle but key point is that the generative model is not encoded by the internal states of the particle being considered: it is not some kind of internal model. Rather, the generative model is a description of the dependence relations that obtain in the dynamics of a system: i.e., such-and-such dynamics entail such-and-such equations of motion and dependencies. FEP-theoretic modelling does not assume that the generative model is encoded in the brain. Arguably, to claim this would indeed constitute a map-territory fallacy. Under the FEP, parts of the system itself (i.e., by internal states) look to an external observer as if, or can be modelled as if, they encode the parameters of a variational model.
However, this raises a question: if it is our scientific model, how can the generative model play a role in inference if only a conditional posterior model is encoded by the system? To claim that a system uses or leverages its generative model is, in some sense, a shorthand expressing the fact that the system leverages its own causal structure, and the data patterns that it elicits through action, to generate patterns of adaptive behaviour—all of which are subject to the regularities that are captured in the generative model. This is a statement of embodied and enactive cognition [50, 13]. The system, in having sensory states, also has access to the causal structure in the generative model.
So, the generative model is our model or map, but it denotes the causal structure of the system being modelled (i.e., the territory). This causal structure—and in particular, the presence of a Markov blanket or particular partition—means that some parts of the system (the internal states) encode a variational density over some other parts of the system (the external states); that is, they “act like a map” or model the statistics of external states.
6.2 Generative models and generative processes
The second ambiguity concerns the generative model and the so-called “generative process”. Some presentations of the FEP sometimes define the generative model as a scientific model of some data-generating generative process; see, for instance, [51, 13]). On this reading, the generative model is our “map” of some real-world process, namely, the generative process—which would then correspond to the metaphorical territory.
In practice, however, in FEP-theoretic modelling, this distinction does not map onto that between a scientific model and real-world system, i.e., onto the distinction between the map and territory. Perhaps confusingly, the generative process does not denote (the structure of) the real data-generating process that we are trying to model. In FEP-theoretic formulations, the generative process is also part of our scientific model.
It should be noted that, practically, the distinction between generative model and generative process only comes into play when we are implementing the dynamics of a specific system in simulations. When we write code to simulate the system of interest, typically, we need to specify (i) the dependencies that hold between the variables of a system, which we call the generative model; and (ii) the equations of motion that govern the time evolution of the external states per se in our simulation. The latter is labelled the ‘generative process’, since it generates the dynamics of external states, which are tracked by internal states. The key thing to note, however, is that generative process is also part of our scientific model—it harnesses what we know about the physics of the situation described by the generative model, in which the thing is embedded and ‘lives’.
7 Conclusion
We argued that, in describing ‘things’ or ‘particles’ as estimators or (deflated) representations of the systems to which they are coupled, the FEP-theoretic apparatus does not commit the map territory fallacy, i.e., it is false that applications the FEP necessarily reify aspects of the metaphorical map (i.e., our scientific model), mistakenly taking them to be part of the territory (i.e., the real-world system that we wish to model); although this remains possible in principle. We have further argued that this allegation itself constitutes a map-territory fallacy fallacy. In distinguishing the generative model and the variational density—and the generative model and the generative process from the real-world system that we aim to model—the technology of FEP-theoretic modelling allows us to construct a map of that part of the territory which looks to an external observer as if or behaves as if it were a map, without committing model reification. The FEP is thus, metaphorically, a map of any possible map whatsoever of, or held by, a physical system. One ought to celebrate this ‘territory-map mapping,’ and the FEP is ultimately, at its core, a principled approach to the formalisation of this mapping.